OpenStudy (anonymous):

Use substitution method to find all solutions of the system of equations y=x^2 y=x+12

hero (hero):

y = y x^2 = x + 12

OpenStudy (anonymous):

how do i get x by it's self if they are unlike terms?

hero (hero):

You don't. You create a quadratic equation instead: x^2 - x - 12 = 0 Then you factor it.

hero (hero):

However, there is a series of steps you can perform to get x by itself. It requires completing the square.

hero (hero):

You could perform the following steps if you wanted to literally get x by itself: x^2 = x + 12 x^2 - x = 12 x^2 - x + 1/4 = 12 + 1/4 (x - 1/2)^2 = 49/4 x - 1/2 = +/- 49/4 x = +/- 7/2 + 1/2

hero (hero):

Which simplifies to x = 8/2 = 4 x = -6/2 = -3

OpenStudy (anonymous):

oh ok thank you I think I understand now

hero (hero):

You still have to find the y values. But not only that, you have to find the right solution pairs.

OpenStudy (anonymous):

for that do I just substitute the variables of x for x to find y?